OpenStudy (anonymous):

Find dy/dx of y=cos(6x+5)

OpenStudy (anonymous):

I know its the chain rule, i already got. Cos(6x+5) * 6

Parth (parthkohli):

That... should be it.

OpenStudy (anonymous):

Using u substitution, but i'm not sure if i did that quite right, Does cos become -sin?

Parth (parthkohli):

Oh yes

Parth (parthkohli):

Wait.

OpenStudy (anonymous):

So -sin(6x+5)*6? Can that simplify more?

Parth (parthkohli):

\[f'(g(x))g'(x) = -\sin(6x + 5)\cdot 6\]Yes, you may use \(-\sin(x) = \sin(-x)\) but you don't need it. So leave it at \(-6\sin(6x + 5)\)

OpenStudy (anonymous):

Think you got your trig messed up there lol.

OpenStudy (anonymous):

Derivative of Sin is cos, and cos is -sin.

Parth (parthkohli):

I guess that's exactly what I did there...?

OpenStudy (anonymous):

Well you said -sin=sin :P

OpenStudy (anonymous):

Also the 6 moves up from with the -sin?

Parth (parthkohli):

No... I said \(-\sin(x) = \sin(-x)\). And yeah, \(-\sin(6x + 5) \cdot 6 = -6\sin(6x+5)\)

OpenStudy (anonymous):

Oh i see, You're right thats a trig identity, but not for derivatives.

OpenStudy (anonymous):

Thanks for the help again : ).

Parth (parthkohli):

:-\ I think \(-\sin(x) = \sin(-x)\) everywhere. You don't need to use that here though. Hmm

OpenStudy (anonymous):

Yeah it probably is, just didnt need it for this.

OpenStudy (anonymous):

You still here, got another question?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

I can't tell if i need the chain rule, or the sum rule here. or both?

OpenStudy (anonymous):

y=(cscx+cotx)^-1

Parth (parthkohli):

So in a way you want to do\[\dfrac{1}{\csc(x) + \cot(x)}\]Right?

OpenStudy (anonymous):

I guess technically thats the same thing yeah.

OpenStudy (anonymous):

Seems to make it more complicated imo.

OpenStudy (anonymous):

Or do i use the product rule now?

OpenStudy (anonymous):

Quotient i mean*

Parth (parthkohli):

I don't know the derivatives of \(\cot(x)\) or \(\csc(x)\) so I'd begin with simplification.\[\dfrac{1}{\frac{\cos(x)}{\sin(x)}+\frac{1}{\sin(x)}} = \dfrac{1}{\frac{1 + \cos(x)}{\sin(x)}} = \dfrac{\sin(x)}{1 + \cos(x)}\]

Parth (parthkohli):

Would that make it simple?

OpenStudy (anonymous):

Well the derv of csc(x)=-cscx*cotx and cot(x)=-cscx^2

Parth (parthkohli):

I don't memorize those :-|

OpenStudy (anonymous):

Those are just basic trig functions. I don't think we can do it like that, might lead to the same answer, but ultimately we need dy/dx.

OpenStudy (anonymous):

I haven't memorized them yet either, need to by tomorrow though :P

Parth (parthkohli):

Yeah, you may do it that way. Never stopped you :-P

OpenStudy (anonymous):

Those are just basic trig functions. I don't think we can do it like that, might lead to the same answer, but ultimately we need dy/dx.

OpenStudy (anonymous):

Well then back to my original question. Do i use the chain rule or how would i go about doing it?

Parth (parthkohli):

You can use the quotient rule.

OpenStudy (anonymous):

Too many options i feel like i'm doing something wrong hmmm

Parth (parthkohli):

How do you want to do it?

OpenStudy (anonymous):

Not sure lol. The correct way?

Parth (parthkohli):

All ways are correct :-P Do you want to continue from \(\dfrac{\sin(x)}{1 + \cos(x)}\)?

OpenStudy (anonymous):

Sure we can. I was doing the quotient rule on my paper.

OpenStudy (anonymous):

Actually wait, is this right so far.

OpenStudy (anonymous):

\[y=\frac{ \csc(x)\cot(x)+\csc^2x }{ (\csc(x)\cot(x))^2 }\]

OpenStudy (anonymous):

Using the quotient rule and trig identities i got there.Maybe could cancel out terms, but not sure.

Parth (parthkohli):

Could you show me the rest of your work? Yes, you could go with cancellation.

OpenStudy (anonymous):

Sure. Well we started with 1/csc(x)cot(x) like you suggested. Then i did the quotient rule. \[(\csc(x)+\cot(x))(0)-(1)(-\csc(x)\cot(x)-\csc^2x) \]

OpenStudy (anonymous):

Which equalled what i just posted above ^^

OpenStudy (anonymous):

Because you divided by (g(x)^2

Parth (parthkohli):

Yeah, good job. The rest?

OpenStudy (anonymous):

I stopped there cause you wanted to see work.

OpenStudy (anonymous):

Will be left with, \[\frac{ \csc^2x }{ \csc(x)\cot(x) }\]

OpenStudy (anonymous):

Final answer? Not really sure to be honest.

OpenStudy (anonymous):

oops actually, the top will just be csc(x) not squared.

OpenStudy (anonymous):

Well whatever thats the right answer, just checked. Thanks for the help.

Parth (parthkohli):

Hmm. That'd give you \(\frac{1}{\cot(x)} = \tan(x)\)